OpenStudy (anonymous):

1. A model is made of a car. The car is 9 feet long, and the model is 6 inches long. What is the ratio of the length of the car to the length of the model? (1 point)

OpenStudy (anonymous):

9:6

OpenStudy (anonymous):

1. A model is made of a car. The car is 9 feet long, and the model is 6 inches long. What is the ratio of the length of the car to the length of the model? (1 point)18 : 1 1 : 18 9 : 6 6 : 9 2. Red and gray bricks were used to build a decorative wall. The H_Geomb_01_01_Proving Triangles Similar_Q2was . There were 175 bricks used in all. How many red bricks were used? (1 point)25 125 50 35 3. A salsa recipe uses green pepper, onion, and tomato in the extended ratio 1 : 3 : 9. How many cups of onion are needed to make 117 cups of salsa? (1 point)3 27 9 81 4. The two rectangles are similar. Which is the correct proportion for corresponding sides? (1 point) 5. You want to draw an enlargement of a design that is printed on a card that is 4 in. by 5 in. You will be drawing this design on a piece of paper that is 8 in. by 11 in. What are the dimensions of the largest complete enlargement you can make? (1 point)1 in. by 10in. 1 in. by in. 8 in. by 4in. 8 in. by 10in. 6. Michele wanted to measure the height of her school’s flagpole. She placed a mirror on the ground 48 feet from the flagpole, then walked backward until she was able to see the top of the pole in the mirror. Her eyes were 5 feet above the ground, and she was 12 feet from the mirror. Using similar triangles, find the height of the flagpole to the nearest tenth of a foot. (1 point)20 ft 38.4 ft 55 ft 25 ft 7. Short Answer Note: Your teacher will grade your responses to questions 7–13 to ensure you receive proper credit for your answers. Note: Remember to show all the steps that you use to solve the problem. You can use the comments field to explain your work. Your teacher will review each step of your response to ensure you receive proper credit for your answer. What is the solution to the proportion? (2 points) 8. Given the proportion a over 8 = 8 over 15, what ratio completes the equivalent proportion a over 8=? (2 points) 9. The polygons are similar, but not necessarily drawn to scale. Find the value of x. (2 points) 10. If , what is the value of A over B? (2 points) 11. Are the two triangles similar? How do you know? triangle diagram (2 points) 12. State whether the triangles are similar. If so, write a similarity statement and the postulate or theorem you used. The diagram is not drawn to scale. lesson 3 quiz question 10 (2 points) 13. Name one location where you might see fractals in everyday life. (1 point)

OpenStudy (anonymous):

#2.5*175/7=125 #3f a salsa recipe uses green pepper, onion, and tomato in the extended ratio 1:3:9, how many cups of onion are needed to make 117 cups of salsa? The answer to this math question is, you would need 27 cups of onion to make 117 cups of salsa. #4 wheres the rest of the numbers in the question? #5 D. 8.5/4 = 2.125 2.125 x 5 = 10.625 or 10 5/8 inch #6.he answer is A. 20 ft ... #7 i need the equation to see it..and answer. #8 a/b = 8/15 means (a/b)(b/8) = (8/15) (b/8), giving a/8 = b/15 #9 i need u to show the graph/ #10-#12...i need to use the graph!

OpenStudy (anonymous):

or atleast see the equasion..and #13... http://www.unc.edu/~jimlee/JohnObrienFractalMusic.htm

OpenStudy (anonymous):

Problem nubmber 4

OpenStudy (anonymous):

Problem #9

OpenStudy (anonymous):

i am not good @ graphs":/ but i'll try

OpenStudy (anonymous):

problem #11

OpenStudy (anonymous):

problem #12

OpenStudy (anonymous):

Can you help me put them into steps please :)

OpenStudy (anonymous):

9ft*12in./1ft.=108in.to 6 inchesor 18 to 1

OpenStudy (anonymous):

Anyone have these answers??